OpenStudy (anonymous):

a rectangular flower bed in front of a building measures 10 1/2 feet bt 1 3/7 feet. what is the total area of the flower bed?

OpenStudy (bahrom7893):

A = (10 1/2) * (1 3/7) = 21/2 * 10/7 = 210/14

OpenStudy (bahrom7893):

A = 15 ft^2 Please click on become a fan if I helped, I really want to get to 100!! Thanks =)